# Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Diverse Terrains

### v2

@@ Abstract
Humanoid robots promise transformative capabilities for industrial and service applications. While recent advances in Reinforcement Learning (RL) yield impressive results in locomotion, manipulation, and navigation, the proposed methods typically require enormous simulation samples to account for real-world variability. This work proposes a novel one-stage training framework—Learn to Teach (L2T)—which unifies teacher and student policy learning. Our approach recycles simulator samples and synchronizes the learning trajectories through shared dynamics, significantly reducing sample complexities and training time while achieving state-of-the-art performance. Furthermore, we validate the RL variant (L2T-RL) through extensive simulations and hardware tests on the Digit robot, demonstrating zero-shot sim-to-real transfer and robust performance over 12+ challenging terrains without depth estimation modules. Experimental videos are available online  [^1] https://lidar-learn-to-teach.github.io.
@@ I Introduction
Mitigation of teacher-student policy divergence: We propose a sample mixing strategy to alleviate the imitation gap between the teacher and student, which traditional privileged learning is unable to address [10]. Both agents will contribute to the replay buffer following a predefined schedule when collecting samples. Mixing samples enables a joint optimization process that mitigates policy divergence while promoting sample efficiency by letting both agents explore Out-of-Distribution (OOD) data.
@@ II Related work
Learning from demonstrations: Learning from demonstrations (LfD) has attracted significant interest in the robot learning field due to the growing abundance of robot data and the popularity of simple yet effective imitation learning (IL) frameworks [15]. LfD has demonstrated impressive results in controlling robot manipulators for tabletop tasks [16]. A recent surge of LfD studies in humanoids and bipeds have shown the promising potential of whole-body control and loco-manipulation [17, 18, 19, 16]. However, supervised learning demands high-quality behavior data, oftentimes through elaborate data collection pipelines [19, 16] and/or needs accurate re-targeting to robot states from datasets with different morphologies , e.g., internet video of human walking. The latter shares growing popularity due to its promising direction towards a generalist robotic control. However, supervised learning can fail to learn reliable whole-body control policies, so they depend on additional model-based balancing controllers [17]. On the other hand, the prevalent IL loss is known to be suboptimal from a pure learning perspective in previous studies [20]. Thus, in this work, we focus on a generic algorithm framework to address the sim2real gap alone, without the interference of possible issues brought up by LfD methods. Furthermore, our proposed framework can be easily extended to the LfD setting, which we leave as a future direction.
@@
Bipedal locomotion over complex terrain: Humanoid robots recently have gained increasing interest due to their applicability and versatility [18, 21, 22, 17, 23, 24, 19, 25, 26, 27, 28], ranging from locomotion [29], to manipulation [30]. Prior bipedal locomotion works [29, 31] have explored the conventional teacher-student learning paradigm in locomotion tasks. However, the training process can take significant samples even with an elaborate training environment design. For instance, in [29], the teacher and the student agent require $1.7e9$ samples. In sharp contrast, we only require $3e8$ samples to train the same Digit robot to walk over challenging terrains for the same control task with even more terrain variations. Concurrent works also incorporate memory structure into the policy architecture [11, 29, 32], or learning from demonstrations collected from various data sources such as human motion data [33] or data generation using model-based methods [30]. In comparison, we design straightforward and intuitive reward functions for bipeds in general terrain settings, offering a simple yet effective solution.
@@ III Methods
We train the teacher with a generic actor-critic method. During training, the teacher interacts with the environment, generates samples, and stores them in a replay buffer [34]. In standard teacher-student frameworks, the teacher’s collected samples are used solely for training the teacher policy $\pi_{t}$ and then discarded. In contrast, our L2T framework co-trains the student with the teacher, reusing the teacher’s samples across all iterations. Fig. 2 illustrates our learning framework, and Algorithm 1 presents the pseudo-code. Specifically, as the teacher interacts with the environment, we record samples $(s,a,r,s^{\prime})$, where $s^{\prime}$ denotes the next observation, the corresponding noisy observations $o$ generated by domain randomization, and $o^{\prime}$, the next observation. In other words, we store $(s,o,a,r,s^{\prime},o^{\prime})$ as training data in the replay buffer. The student updates its policy at each iteration by sampling mini-batches from the replay buffer, but its policy relies solely on the collected noisy data. This joint training procedure greatly improves sample efficiency as both agents learn together, without the need for a separate student training stage as that used in the traditional set-up.
@@
To systematically blend teacher and student experiences, we define a mixture coefficient $\alpha_{\text{mix}}$. At each time step, the action $a$ is determined by a probabilistic mixture of the teacher’s policy $\pi_{t}$ and the student’s policy $\pi_{s}$. Specifically, the action selection is defined as follows:
@@
Additionally, $\alpha_{\text{mix}}$ is scheduled linearly from 0 to a predefined constant over the course of training, which is $0.2$ in our implementation. This formulation ensures that, initially, the teacher’s guidance dominates the action selection, but as training progresses, the student’s policy increasingly influences the learning. The scheduling of $\alpha_{\text{mix}}$ helps balance the contributions of both policies and guarantees the stability of the training process. By combining these strategies, our framework leverages both the teacher’s guidance and the student’s explorative capabilities to achieve more robust learning outcomes.
@@
where the optimal $p_{t}$ represents the teacher policy in the next iteration. Any policy improvement scheme can be fit into the L2T framework. This formulation encompasses a range of policy gradient methods [35], such as Proximal Policy Optimization (PPO) [36] and Soft Actor-Critic (SAC) [37]. In our experiments on the Digit robot, we employ a PPO-style update. However, our framework can be easily extended to various learning methods, including imitation learning methods, or Inverse Reinforcement Learning (IRL) methods such as in [20].
@@
$\min_{p_{s}}\;L_{IL}=\mathbb{E}_{s,\,o\sim D}\left\|p_{s}(\cdot|o)-\pi_{t}(\cdot|s)\right\|_{2},$ (3)
@@
$\min_{p_{s}}\;L_{KL}=\mathbb{E}_{s,\,\sim D}\;\text{KL}\left(p_{s}(\cdot|o)\;\|\;\pi_{t}(\cdot|s)\right),$ (4)
@@
$\min_{p_{s}}L_{\text{Asym}}=-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{s}(\cdot|o)\rangle+\text{KL}(p_{s}\|\pi_{t}).$ (5)
We denote the general loss function for student agents as $L_{s}$. In our application on the Digit robot, we observed that using the $L_{2}$ imitation loss yields the best performance, while the addition of $L_{\text{Asym}}$ does not affect the overall performance by a large margin. We conjecture that the $L_{2}$ loss allows the student policy to have a slightly higher exploration capability as we observe that $L_{\text{Asym}}+L_{2}$ will reach a training plateau that is inferior in performance than using $L_{2}$ alone.
@@
1:Input: initial teacher policy $\pi_{t}^{0}$, student policy $\pi_{s}^{0}$, and step size sequences $\{\beta^{k}_{t}\}$ and $\{\beta^{k}_{s}\}$.
2:for $k=0$ to $K$ do
3:  Sample a mini-batch $D_{k}$ from replay buffer $D$.
4:  Update the teacher critic by estimating:
$Q^{\pi_{t}^{k+1}}(s,a),V^{\pi_{t}^{k+1}}(s).$
5:  Update the teacher policy:
$\pi^{k+1}_{t}=\arg\min_{p_{t}}-\beta^{k}_{t}\langle Q^{\pi_{t}^{k}}(s,\cdot),p_{t}(\cdot|s)\rangle+\text{KL}(\pi_{t}^{k}\|p_{t}).$
6:  Update the student policy:
$\pi^{k+1}_{s}=\arg\min_{\pi_{s}}L_{s}(\pi_{s}^{k}).$
7:  Roll out to collect new samples $D^{\prime}$ according to the scheduling in Eq. 1.
8:  Update the replay buffer: $D\leftarrow D\cup D^{\prime}$.
9:end for
@@ IV Environment Design
The Digit robot is a bipedal walking robot with $30$ degrees of freedom, which includes $20$ actuated joints with $4$ per arm and $6$ per leg. All joints are revolute joints except for the shin and heel joints, which are spring-based. Notably, the Digit robot features three closed kinematic chains per leg. Two of these chains involve motors controlling the foot, assisted by additional rods, while the third chain is responsible for controlling the heel via a rod extending from the hip. This leg design makes it a significantly challenging task for RL algorithms due to the high-dimensional action space and the complex dynamics of the robot.
We highlight a significant portion of our work is to reconstruct a faithful Universal Scene Description (USD) model of the robot in IsaacLab [9], although this is not claimed as an algorithmic contribution. As a result, we build a velocity-tracking RL task with accurate dynamics w.r.t the robot hardware, facilitating the sim2real transfer.
@@ IV-A Observation space
The observation space (see table I) is constructed using data provided by the robot’s sensors, including base linear velocity, base angular velocity, joint positions, and joint velocities. Additionally, we include the commanded velocity that the robot will receive from an external controller during execution and the computed projected gravity and gait phase based on the IMU data and the robot execution time. Finally, the actions in a previous time step are also incorporated into the observation space, which allows us to learn a history-dependent policy using recurrent neural nets. We model measurement noise as $o_{t}=s_{t}+\alpha\epsilon$ where $\alpha$ is the scale, and $\epsilon$ is either Gaussian or uniform noise. We add this noise to the student’s observation space to mimic the hardware sensors while keeping the teacher’s observation noise-free, except for the ones that incorporate the student’s observations in order to alleviate the imitation gap, which is considered a common practice. Additionally, privileged information (lower half of Table I) is provided for the teacher for easier training.
@@
Base lin. vel. 3
Base ang. vel. 3
Joint position 48
Joint velocity 48
@@ IV-C Reward functions
Joint deviation hip -5.0 Feet air time 0.125
Joint deviation arms -0.3 Joint deviation toes -0.1
@@
$h_{\text{foot\_traj}}$ is the actual foot height and $h_{\text{foot\_traj\_target}}$ is the target foot height. We adjust the desired foot height based on the current CoM position to adapt to uneven terrains. Specifically, instead of tracking the absolute height of the foot, we track the relative distance of the foot and the CoM position to compensate for the terrain height, which is hard to obtain at run time. This reward seems to be the most important reward term, without which we are unable to train an agent robust to challenging terrains.
@@ IV-D Domain randomization and curriculum
Additionally, we adopt the curriculum training setup implemented by IsaacLab. We utilize the existing terrain map, which includes seven different terrains: flat ground, slopes, stepping stones, and pyramid stairs up and down. Each environment, except for flat, has a difficulty parameter from 0 to $1$, where 0 stands for the easiest and $1$ the hardest. As an example, for the slope environment, the degree of the slope ranges from 0 to $40$. The environment is configured with a curriculum such that the agents’ performance will increase if they improve their tracking of the generated velocity command and decrease if they fail to do so.
TABLE IV: Neural Network Parameters
Parameter Value
Number of GPUs 1 RTX 4080
Isaac Lab Decimation 20
Simulation Time Step (dt) 0.001 s
Number of Environments 4096
Learning Epochs 5
Steps per Environment 12
Batch Size 256
Episode Length 1000
Discount Factor ($\gamma$) 0.99
Generalized Advantage Estimation ($\lambda$) 0.95
Entropy Coefficient 0.001
Clipping Parameter 0.2
Optimizer Adam
Learning Rate 5e-4
Learning Rate Schedule Linear Decay
Total Time Steps (Samples) 5e8
Number of Iterations (approx) 6100
Figure 3: Training curve against baselines. (a) We plot the difficulty of the terrain the agent can successfully traverse during training. (b) We train and evaluate each method on the flat ground and plot the velocity tracking of each agent.
@@ V-A Rough terrain difficulty progression
First, we show the progression of terrain difficulty (curriculum progression) throughout training, representing the overall task completion, i.e., the level of terrain difficulty the agent can walk over with the CoM velocity maintained within a specific range from the velocity command. This metric is calculated as the average difficulty level across the $6$ terrain setups mentioned earlier. In essence, faster learning corresponds to a steeper curve in terrain difficulty progression. Fig. 3(a) illustrates the terrain progression during training. The purple curve illustrates the training curve of L2T, reaching a maximum of $5.72$ level of difficulty. Compared to a two-stage traditional teacher-student learning (TSL) paradigm, which is represented by two curves: the teal curve represents the TSL teacher agent, and the blue curve represents the TSL student agent. Notice that in TSL, the student agent can only be trained after the teacher. The TSL teacher has an MLP-based policy, while the student is LSTM-based, sharing the same network architecture and hyperparameter as in L2T. It is noteworthy to mention that we can only obtain a reasonable training result by using data aggregation (DAGGER) [38], i.e., when training the student, periodically use the teacher to predict the action. Additionally, we include two RL methods for learning with student observation space using vanilla PPO, and recurrent PPO. It turns out that these two baselines perform significantly inferior. Since IsaacLab cannot be evaluated during training, the L2T curves represent the joint training progression between the L2T teacher and the student. One might think that the true performance of L2T students can be much worse than what the curve shows. However, as we observe that the L2T curve eventually goes higher than the TSL teacher, we argue that this curve is rather the performance lower bound for the L2T student.
V-B Flat terrain velocity tracking error
Next, we investigate the agent velocity tracking performance. Unfortunately velocity tracking is not a valid metric on general terrain due to the variability of the terrain difficulty. For example, the same agent might exhibit significantly different velocity tracking abilities either on slopes or stairs. To make a fair comparison, we re-train separate agents and collect the results on flat terrain, as it presents the most straightforward task for locomotion. We compare the velocity tracking between the CoM velocity and the commanded velocity on the $xy$ plane and the yaw angle. The results are plotted in Fig. 3(b). The L2T agent and the student from the traditional teacher-student learning (TSL) paradigm are comparable, with L2T better at yaw tracking and worse at tracking in the $xy$ plane. This shows that on the flat surface, L2T performs similarly to traditional teacher-student training paradigms but saves $50\%$ of the total samples. It is noteworthy to mention that we use the same network architecture and hyperparameters for both training frameworks. In Fig. 5(b)-(c), we also plot the recorded robot CoM sagittal velocity versus the commanded velocity and the robot’s CoM trajectory in the horizontal plane.
Figure 4: Left: in conventional TSL, the teacher supervises student. Right: L2T co-trains both agents together, where the student provides feedback to the teacher through mixed samples.
Figure 5: (a) Ablation study on the mixture coefficient $\alpha_{\rm mix}$. (b) CoM velocity versus commanded velocity. (c) xyxy trajectory with $x$ direction command.
Figure 6: (a) we perturb the robot with a harness. (b) we test the step-in-place motion on a crate of gravel. We let the robot walk over (c) a slippery terrain with (left) the Agility company controller and (right) our L2T controller. (d) rocky terrain with slopes. (e) elevated platforms. (f) hill with gravel.
V-C Mitigation of policy divergence

### v3

# Learn to Teach: Sample-Efficient Privileged Learning for Humanoid Locomotion over Real-World Uneven Terrain

@@ Abstract
Humanoid robots promise transformative capabilities for industrial and service applications. While recent advances in Reinforcement Learning (RL) yield impressive results in locomotion, manipulation, and navigation, the proposed methods typically require enormous simulation samples to account for real-world variability. This work proposes a novel one-stage training framework—Learn to Teach (L2T)—which unifies teacher and student policy learning. Our approach recycles simulator samples and synchronizes the learning trajectories through shared dynamics, significantly reducing sample complexities and training time while achieving state-of-the-art performance. Furthermore, we validate the RL variant (L2T-RL) through extensive simulations and hardware tests on the Digit robot, demonstrating zero-shot sim-to-real transfer and robust performance over 12+ diverse terrains without depth estimation modules. Experimental videos are available at https://lidar-learn-to-teach.github.io.
Index Terms:
Reinforcement Learning, Humanoid and Bipedal Locomotion, Sim2Real, Sample Efficiency
@@ I Introduction
Mitigation of teacher-student imitation gap: We propose a sample mixing strategy to alleviate the imitation gap between the teacher and student, which traditional privileged learning is unable to address [10]. Both agents will contribute to the replay buffer following a predefined schedule when collecting samples. Mixing samples enables a joint optimization process that mitigates the imitation gap while promoting sample efficiency by letting both agents explore Out-of-Distribution (OOD) data.
@@ II Related work
Learning from demonstrations: Learning from demonstrations (LfD) has attracted significant interest in the robot learning field due to the growing abundance of robot data and the popularity of simple yet effective imitation learning (IL) frameworks [15]. LfD has demonstrated impressive results in controlling robot manipulators for tabletop tasks [16]. A recent surge of LfD studies in humanoids and bipeds have shown the promising potential of whole-body control and loco-manipulation [17, 18, 19, 16]. However, supervised learning demands high-quality behavior data, oftentimes through elaborate data collection pipelines [19, 16] and/or needs accurate re-targeting to robot states from datasets with different morphologies On the other hand, the prevalent IL loss is known to be suboptimal from a learning perspective [20]. Thus, in this work, we focus on a generic algorithm framework to address the sim2real gap alone, without the interference of possible issues brought up by LfD methods. Furthermore, our proposed framework can be easily extended to the LfD setting, which we leave as a future direction.
@@
Bipedal locomotion over complex terrain: Humanoid robots recently have gained increasing interest due to their applicability and versatility [18, 21, 22, 17, 23, 24, 19, 25], ranging from locomotion [26], to manipulation [27]. Prior bipedal locomotion works [26, 28] have explored the conventional teacher-student learning paradigm in locomotion tasks. However, the training process can take significant samples even with an elaborate training environment design. Concurrent works also incorporate memory structure into the policy architecture [11, 26, 29], or learning from demonstrations collected from various data sources such as human motion [30] or generation using model-based methods [27]. In comparison, we design straightforward and intuitive reward functions for bipeds in general terrain settings, offering a simple yet effective solution.
@@ III Methods
We train the teacher with a generic actor-critic method. During training, the teacher interacts with the environment, generates samples, and stores them in a replay buffer [31]. In standard teacher-student frameworks, the teacher’s collected samples are used solely for training the teacher policy $\pi_{t}$ and then discarded. In contrast, our L2T framework co-trains the student with the teacher, reusing the teacher’s samples across all iterations. Fig. 2 illustrates our learning framework, and Algorithm 1 presents the pseudo-code. We employ an MLP for the teacher because, given privileged access to the full state (depth scans, root pose, terrain profile, etc.), each observation is already fully descriptive and Markovian so that there is no need to model temporal dependencies via recurrence. We also notice that this architecture has been widely adopted by concurrent teacher-student learning paradigms [26]. Specifically, as the teacher interacts with the environment, we record samples $(s,a,r,s^{\prime})$, where $s^{\prime}$ denotes the next observation, the corresponding noisy observations $o$ generated by domain randomization, and $o^{\prime}$, the next observation. In other words, we store $(s,o,a,r,s^{\prime},o^{\prime})$ as training data in the replay buffer. The student updates its policy at each iteration by sampling mini-batches from the replay buffer, but its policy relies solely on the collected noisy data. This joint training procedure greatly improves sample efficiency as both agents learn together, without the need for a separate stage as used in the traditional set-up.
@@
To systematically blend teacher and student experiences, we define a mixture coefficient $\alpha_{\text{mix}}$. At each time step, the action $a$ is determined by a probabilistic mixture of the teacher’s policy $\pi_{t}$ and the student’s policy $\pi_{s}$. As such, we collect trajectories $s_{1},a_{1},s_{2},a_{2},s_{3},a_{3}$ where $a_{i}$ could either come from the teacher or the student. This sample mixing mechanism ensures that information from the student agent back propagates to the teacher. Specifically, the action selection is defined as follows:
@@
Additionally, $\alpha_{\text{mix}}$ is scheduled linearly from 0 to a predefined constant over the course of training, which is $0.2$ in our implementation. This formulation ensures that, initially, the teacher’s guidance dominates the action selection, but as training progresses, the student’s policy increasingly influences the learning. The scheduling of $\alpha_{\text{mix}}$ helps balance the contributions of both policies and guarantees stable training. By combining these strategies, our framework leverages both the teacher’s guidance and the student’s explorative capabilities to achieve more robust learning outcomes. This is in steep contrast to CTS, where the teacher encoder and the student encoder trains on independent datasets.
@@
where the optimal $p_{t}$ represents the teacher policy in the next iteration. Any policy improvement scheme can be fit into the L2T framework. This formulation encompasses a range of policy gradient methods [32], such as Proximal Policy Optimization (PPO) [33] and Soft Actor-Critic (SAC) [34]. In our experiments on the Digit robot, we employ a PPO-style update. However, our framework can be easily extended to various learning methods, including imitation learning methods, or Inverse Reinforcement Learning (IRL) methods such as in [20]. In practice, we pool samples from both agents to update the teacher policy using PPO. If we can maintain $\pi_{s}\approx\pi_{t}$, any student action $a_{s}\sim\pi_{s}(\cdot|s)$ satisfies $\pi_{t}(a_{s}|s)>0$, allowing us to view these student-generated transitions as valid (albeit lower-probability) samples from the teacher’s distribution.
@@
$\min_{p_{s}}\;L_{\text{IL}}=\mathbb{E}_{s,\,o\sim D}\left\|p_{s}(\cdot|o)-\pi_{t}(\cdot|s)\right\|_{2},$ (3)
@@
$\min_{p_{s}}\;L_{KL}=\mathbb{E}_{s,\,o\sim D}\;\text{KL}\left(p_{s}(\cdot|o)\;\|\;\pi_{t}(\cdot|s)\right),$ (4)
@@
$\min_{p_{s}}L_{\text{Asym}}=-\beta\langle Q^{\pi_{t}}(s,\cdot),p_{s}(\cdot|o)\rangle+\text{KL}(p_{s}\|\pi_{s}).$ (5)
We denote the general loss function for student agents as $L_{s}$. In our application on the Digit robot, we observed that using the $L_{\text{IL}}$ imitation loss yields the best performance, while the addition of $L_{\text{Asym}}$ does not affect the overall performance by a large margin. We conjecture that the $L_{\text{IL}}$ loss allows the student policy to have a slightly higher exploration capability as we observe that $L_{\text{Asym}}+L_{\text{IL}}$ will reach a training plateau that is inferior in performance than using $L_{\text{IL}}$ alone.
@@
0: initial teacher policy $\pi_{t}^{0}$, student policy $\pi_{s}^{0}$, and step size sequences $\{\beta^{k}_{t}\}$ and $\{\beta^{k}_{s}\}$
1:for $k=0$ to $K$ do
2:  Sample a mini-batch $D_{k}$ from the replay buffer $D$
3:  Update the teacher critic $Q^{\pi_{t}^{k+1}}(s,a),V^{\pi_{t}^{k+1}}(s)$
4:  Update the teacher policy:
$\pi^{k+1}_{t}=\arg\min_{p_{t}}\left[-\beta^{k}_{t}\langle Q^{\pi_{t}^{k}}(s,\cdot),p_{t}(\cdot|s)\rangle+\mathrm{KL}(\pi_{t}^{k}\|p_{t})\right]$
5:  Update the student policy:
$\pi^{k+1}_{s}=\arg\min_{\pi_{s}}L_{s}(\pi_{s}^{k})$
6:  Roll out to collect new samples $D^{\prime}$ according to the scheduling in Eq. 1
7:  Update the replay buffer: $D\leftarrow D\cup D^{\prime}$
8:end for
@@ IV Environment Design
The Digit robot is a bipedal walking robot with $30$ degrees of freedom, which includes $20$ actuated joints with $4$ per arm and $6$ per leg. All joints are revolute joints except for the shin and heel joints, which are spring-based. Notably, the Digit robot features three closed kinematic chains per leg. Two of these chains involve motors controlling the foot, assisted by additional rods, while the third chain is responsible for controlling the heel via a rod extending from the hip. This leg design makes it a significantly challenging task for RL algorithms due to the high-dimensional action space and the complex dynamics of the robot. We highlight a significant portion of our work is to reconstruct a faithful Universal Scene Description (USD) model of the robot in IsaacLab [9], although this is not claimed as an algorithmic contribution. As a result, we build a velocity-tracking RL task with accurate dynamics w.r.t the robot hardware.
@@ IV-A Observation space
The observation space (see table I) is constructed using data provided by the robot’s sensors, including base linear velocity, base angular velocity, joint positions, and joint velocities. Additionally, we include the commanded velocity that the robot will receive from an external controller during execution, the computed projected gravity based on the IMU data, and desired gait phase based on the robot execution time. Finally, the actions in a previous time step are also incorporated into the observation space, which allows us to learn a history-dependent policy using recurrent neural nets. We model measurement noise as $o_{t}=s_{t}+\alpha\epsilon$ where $\alpha$ is the scale, and $\epsilon$ is either Gaussian or uniform noise. We add this noise to the student’s observation space to mimic the hardware sensors while keeping the teacher’s observation noise-free, except for the ones that incorporate the student’s observations in order to alleviate the imitation gap, which is considered a common practice. Additionally, privileged information (lower half of Table I) is provided for the teacher for easier training. While our locomotion experiments randomize only the base mass, we include all environment parameters as privileged information, bundling them simplifies our environment API for future extensions. Note that the base linear velocity is given by the low-level software APIs provided by Agility Robotics.
@@ IV-C Reward functions
Joint deviation hip -5.0 Joint deviation toes -0.1
Joint deviation arms -0.3
@@
$h_{\text{foot\_traj}}$ is the actual foot height and $h_{\text{foot\_traj\_target}}$ is the target foot height. We adjust the desired foot height based on the current CoM position to adapt to uneven terrains. Specifically, instead of tracking the absolute height of the foot, we track the relative distance of the foot and the CoM position to compensate for the terrain height, which is hard to obtain at run time. This reward produces significantly different motions than rewarding foot clearance alone and it is one of the most important reward term, without which we are unable to train an agent robust to uneven terrains. Note that, this reward is different from the foot clearance reward as the latter rewards the foot to reach a desired height relative to floor. On flat ground, the desired height is consistent with the peak of the reference foot trajectory.
@@ IV-D Domain randomization and curriculum
Additionally, we adopt the curriculum training setup implemented by IsaacLab. We utilize the existing terrain map, which includes seven different terrains: flat ground, slopes, stepping stones, and pyramid stairs up and down.
Figure 3: Training curve against baselines. Left: We plot the difficulty of the terrain the agent can successfully traverse during training. Middle: we show the student training curve in TSL after obtaining a teacher. Right: average success rate of the student agents across $4096$ training environments with randomly sampled velocity commands.
@@ V-A Rough terrain difficulty progression
First, we show the progression of terrain difficulty (curriculum progression) throughout training, representing the overall task completion, i.e., the level of terrain difficulty the agent can walk over with the CoM velocity maintained within a specific range from the velocity command. This metric is calculated as the average difficulty level across the $6$ terrain setups mentioned earlier. In essence, faster learning corresponds to a steeper curve in terrain difficulty progression. Fig. 3 illustrates the terrain progression during training. We represent the teacher policy with a solid blue line and the student policy with a dashed blue line. We compare our approach against three baseline methods based on the asymmetric learning framework: CTS [11], ROA [35], and DWL [36]. In addition, we include comparisons with the conventional teacher-student learning (TSL) paradigm and a vanilla recurrent PPO trained directly in the student’s environment. Notice that in TSL, the student agent can only be trained after the teacher. The TSL teacher has an MLP-based policy, while the student is LSTM-based, sharing the same network architecture and hyperparameter as in L2T. It is noteworthy to mention that we can only obtain a reasonable training result by using data aggregation (DAgger) [37], i.e., when training the student, periodically use the teacher to predict the action.
Discussion: First, Fig. 3 shows that L2T significantly outperforms asymmetric learning baselines, including CTS, ROA, and DWL. We are unable to reproduce competitive performance with DWL due to missing implementation details. However, as DWL follows a similar framework to ROA, learning an encoder to reconstruct privileged states, we believe its potential for improvement over ROA is limited. Since these (single actor) methods lack a separate teacher policy, the student must discover high-reward states with a suboptimal student encoder in order to learn from them, which leads to poor sample efficiency. Furthermore, we observe a substantial imitation gap between the teacher and student in CTS. This arises because CTS processes teacher and student samples independently, relying solely on a reconstruction loss to align the two encoders. In contrast, L2T mixes samples from both agents, allowing the teacher to observe and evaluate the student’s behavior. From a data-driven viewpoint, CTS distills knowledge using two agents trained on disjoint datasets, while L2T actively injects out-of-distribution student trajectories into the teacher’s training, enabling more effective guidance of the student policy.
Second, compared to the conventional teacher-student learning (TSL) paradigm, we find that L2T enables the student to achieve comparable performance without requiring a separate teacher pretraining stage. That is, the time spent training the teacher in TSL is sufficient for jointly training both teacher and student in L2T, making it a drop-in replacement. While we do not claim that L2T outperforms TSL in final policy quality, we save equivalently $50\%$ training time, or roughly $12$ hours of GPU time, a significant improvement in reducing development cycles. Ultimately, both approaches are upper-bounded by the performance of the teacher policy.
(a) Imitation gap ablation.
(b) Loss function ablation.
Figure 4: Ablation studies on sample mixing and loss functions.
V-B Mitigation of imitation gap